OpenStudy (anonymous):

I need to find the acute angle on a right angled triangle with an opposite of 18 and an adjacent of 70 please

OpenStudy (gorv):

welcome to os :)

OpenStudy (gorv):

can u be more specific ??

OpenStudy (bradely):

Use tan theta = opp/adj = 18/70 =0.257 theta = tan^-1(0.257) =14.42 Source: http://www.mathskey.com/question2answer/